Suppose that business travelers and vacationers have the following demand for airline tickets from New York to Boston:


Price Quantity Demanded (Business travelers) Quantity Demanded (Vacationers)


$150 2,100 tickets 1,000 tickets


200 2,000 800


250 1,900 600


300 1,800 400


a) As the price of tickets rises from $200 to $250, what is the price elasticity of demand for (i) business travelers and (ii) vacationers? (Use the midpoint method in your calculations.


b) Why might vacationers have a different elasticity than business travelers?


1
Expert's answer
2021-12-22T13:30:06-0500

1) Price elasticity of demand for business travellers.

Formula.

Percent change in quantity demanded/percent change in price.

Percent change in quantity demanded "= \frac{q_2-q_1}{q_1}"

"=\frac{2000-1900}{2000} \\

=\frac{100}{2000} \\

=0.05"

Percent change in price "= \frac{p_2-p_1}{p_1}"

"=\frac{200-250}{200} \\

=\frac{-50}{200} \\

=-0.25"

Therefore.

"\frac{0.05}{-0.25}=-0.2"

Therefore price elasticity of demand for business travellers=-0.2

2) For vacationers.

Percent change in quantity demanded

"= \frac{800-600}{800} \\

=\frac{200}{800} \\

=0.25 \\

\frac{0.25}{-0.25} =-1"

Therefore price elasticity of demand for vacationers=-1


b)

Vacationers are casual with their plan whereas business travelers are traveling with business purpose thus their elasticity of demand for tickets is less affected with change in price.